i need to know how to work this problem
2(1/4+1/5)+2 please help

Found 2 solutions by stanbon, ReadingBoosters:
Answer by stanbon(75887)

2(1/4+1/5)+2
-----
2[(5+4)/20]+2
----
= 2(9/20)+2
----
= (18/20) + (40/20)
--------
= (58/20)
-----
= 2 9/10
====================
Cheers,
Stan H.

Answer by ReadingBoosters(3246)

Add 1/4 + 1/5 by establishing common denominator
1/4(5/5) + 1/5(4/4)
5/20 + 4/20 = 9/20
2(9/20) + 2
18/20 + 2
9/10 + 2 = 2 9/10 OR
Convert 2 into fraction with denominator of 10
10*2 = 20
9/10 + 20/10 = 29/10
Convert to mixed fraction
2 9/10